OpenStudy (reneeshumpert):

x - y = 4 x + y = 8 The x-coordinate of the solution to the system shown is _____. My work x-y=4 - x+y=8 --------- 0 - 2y = 4 1. Divide two on both sides y= -2 x-2=4 Add two to both sides x=6 Second equation 6+ -2= 8 So my answer is 6 ?

TheSmartOne (thesmartone):

\(\sf 4-8 = -4\)

OpenStudy (anonymous):

x−y=4;−x+y=8 Step: Solve x−y=4 for x: x−y+y=4+y(Add y to both sides) x=y+4 Step: Substitute (y+4) for x in −x+y=8: −x+y=8 −(y+4)+y=8 −4=8(Simplify both sides of the equation) −4+4=8+4(Add 4 to both sides) ?????

OpenStudy (reneeshumpert):

2 4 6 12 My answer options

TheSmartOne (thesmartone):

So restart from here - 2y = -4 y= 2

OpenStudy (anonymous):

Look at the steps and solve the last step

TheSmartOne (thesmartone):

You did that error but you came out with the same correct answer... interesting coincidence lol

OpenStudy (reneeshumpert):

I knew I did something wrong somewhere because at first the equations werent coming equal, but so the answer is 6?

OpenStudy (anonymous):

My answer is 12 -_-

OpenStudy (reneeshumpert):

I dont think " x " could be 12 lol

TheSmartOne (thesmartone):

It is not 12. y= 2 So x-y=4 x-2=4 x=6 and also x+y=8 x+2=8 x=6 So yes 6 is correct :)

OpenStudy (reneeshumpert):

Yes ! Well I think I got this section of my work under wraps ill be back with my next lesson in about an hour thank you for the help

OpenStudy (anonymous):

Well sorry about the wrong answer :/

OpenStudy (reneeshumpert):

its cool